# Non-paraxial effects on laser-qubit interactions

### v1

@@ Appendix A Squeeze and displacement transformation
with quadrature transformations $\hat{D}^{\dagger}(\alpha\hat{\sigma}_{z})\hat{a}\hat{D}(\alpha\hat{\sigma}_{z})=\hat{a}-\alpha\hat{\sigma}_{z}$ and $\hat{D}^{\dagger}(\alpha\hat{\sigma}_{z})\hat{a}^{\dagger}\hat{D}(\alpha\hat{\sigma}_{z})=\hat{a}^{\dagger}-\alpha^{*}\hat{\sigma}_{z}$. The Hamiltonian becomes
$\tilde{H}=$ $\tilde{\omega}\left(\hat{a}^{\dagger}\hat{a}+1/2-\alpha\hat{\sigma}_{z}(\hat{a}+\hat{a}^{\dagger})+\alpha^{2}\right)+\delta^{(0)}\hat{\sigma}_{z}+\delta^{(1)}e^{-{\xi\hat{\sigma}_{z}}}x_{0}(\hat{a}+\hat{a}^{\dagger})\hat{\sigma}_{z}$
$-2\delta^{(1)}e^{-{\xi\hat{\sigma}_{z}}}x_{0}\alpha+\hat{D}^{\dagger}(\alpha\hat{\sigma}_{z})\hat{S}^{\dagger}(\xi\hat{\sigma}_{z})\hat{\Omega}\hat{\sigma}_{x}\hat{S}(\xi\hat{\sigma}_{z})\hat{D}(\alpha\hat{\sigma}_{z}),$ (31)
@@
when $\alpha$ is real. The linear motional terms cancel out after choosing the displacement parameter $\alpha=\delta^{(1)}l_{\text{ho}}/\omega$ to be a dimensionless measure of the Stark shift gradient. The Hamiltonian simplifies to
$\tilde{H}=\tilde{\omega}(\hat{a}^{\dagger}\hat{a}+1/2)-\tilde{\omega}\alpha^{2}+\delta^{(0)}\hat{\sigma}_{z}+\hat{D}^{\dagger}(\alpha\hat{\sigma}_{z})\hat{S}^{\dagger}(\xi\hat{\sigma}_{z})\hat{\Omega}\hat{\sigma}_{x}\hat{S}(\xi\hat{\sigma}_{z})\hat{D}(\alpha\hat{\sigma}_{z}).$ (32)
@@
We disregard energy offsets that do not contribute to qubit-motion coupling, like $\delta^{(0)}\hat{\sigma}_{z}$, which can be fixed by the laser frequency. The Hamiltonian is
$\tilde{H}=\tilde{\omega}(\hat{a}^{\dagger}\hat{a}+1/2)+\hat{D}^{\dagger}(\alpha\hat{\sigma}_{z})\hat{S}^{\dagger}(\xi\hat{\sigma}_{z})\hat{\Omega}\hat{\sigma}_{x}\hat{S}(\xi\hat{\sigma}_{z})\hat{D}(\alpha\hat{\sigma}_{z}).$ (33)
@@
The first term on the right hand side describes the ion’s motion as a modified harmonic oscillator with frequency $\tilde{\omega}$. The second term describes the atom-laser coupling acted on by qubit-dependent squeeze and displacement operators. We expand the operator $\hat{\Omega}=\Omega^{(0)}+\Omega^{(2)}\hat{x}^{2}$ and first consider the Rabi frequency amplitude $\Omega^{(0)}$. We utilize the Baker-Hausdorff lemma to determine the non-commutating corrections to the Rabi frequency amplitude. The first two corrections proportional to $\hat{n}$ (i.e. disregarding processes that change the Fock state) in order of magnitude are
$\hat{D}^{\dagger}(\alpha\hat{\sigma}_{z})\hat{S}^{\dagger}(\xi\hat{\sigma}_{z})\left(\Omega^{(0)}\hat{\sigma}_{x}\right)\hat{S}(\xi\hat{\sigma}_{z})\hat{D}(\alpha\hat{\sigma}_{z})=\Omega^{(0)}\left(1+2\alpha^{2}(2\hat{n}+1)+\frac{\xi\alpha}{2}(2\hat{n}+1)+\order{\hat{n}^{2}}\right)\hat{\sigma}_{x}$ (34)
The second correction proportional to $\xi\alpha/2$ is approximately two orders of magnitude smaller than the first (as $\xi\ll\alpha$ for our parameters) and can be neglected, along with higher order terms such as $\xi^{2}$, $\alpha^{3}$ etc. that are linear in $\hat{n}$.
@@
We then consider the Rabi frequency curvature $\Omega^{(2)}\hat{x}^{2}$, which expanded into creation and annihilation operators is $\Omega^{(2)}\hat{x}^{2}\hat{\sigma}_{x}=\Omega^{(2)}l_{\text{ho}}^{2}(2\hat{n}+1+\hat{a}^{\dagger}\hat{a}^{\dagger}+\hat{a}\hat{a})\hat{\sigma}_{x}$. There are additional non-commutating motional terms with the squeeze and displacement operators. To leading order in $\hat{n}$ these are off-resonant, so we find
@@
$\hat{D}^{\dagger}(\alpha\hat{\sigma}_{z})\hat{S}^{\dagger}(\xi\hat{\sigma}_{z})\left(\Omega^{(2)}\hat{x}^{2}\hat{\sigma}_{x}\right)\hat{S}(\xi\hat{\sigma}_{z})\hat{D}(\alpha\hat{\sigma}_{z})$ $=\Omega^{(2)}l_{\text{ho}}^{2}(2\hat{n}+1)\left(1+2\alpha^{2}(2\hat{n}+1)+\frac{\xi\alpha}{2}(2\hat{n}+1)+\order{\hat{n}^{2}}\right)\hat{\sigma}_{x}$ (35)
$=\Omega^{(2)}l_{\text{ho}}^{2}\left((2\hat{n}+1)+8\alpha^{2}\hat{n}+\order{\hat{n}^{2}}\right)\hat{\sigma}_{x}.$ (36)
The first correction $8\alpha^{2}\hat{n}$ is suppressed by $\sim 10^{6}$ so is negligible. The exclusion of off-resonant motional terms in our treatment becomes noticable when $\bar{n}\ll 1$ or $\Omega\gg\omega$, as seen in Fig. 5 compared to the numeric simulations. Grouping together different parts the Hamiltonian is
$\tilde{H}=\omega(\hat{a}^{\dagger}\hat{a}+1/2)+\Omega^{(0)}\hat{\sigma}_{x}+2\xi\omega(\hat{n}+1/2)\hat{\sigma}_{z}+\left(2\Omega^{(0)}\alpha^{2}(2\hat{n}+1)+\Omega^{(2)}l_{\text{ho}}^{2}(2\hat{n}+1)\right)\hat{\sigma}_{x},$ (37)
written in physical parameters as
@@
$\tilde{H}=\omega(\hat{a}^{\dagger}\hat{a}+1/2)+\Omega^{(0)}\hat{\sigma}_{x}+2\delta^{(2)}l_{\text{ho}}^{2}(\hat{n}+1/2)\hat{\sigma}_{z}+\left((2\Omega^{(0)}(\delta^{(1)})^{2}l_{\text{ho}}^{2}/\omega^{2})(2\hat{n}+1)+\Omega^{(2)}l_{\text{ho}}^{2}(2\hat{n}+1)\right)\hat{\sigma}_{x}.$ (38)
In the text we write this Hamiltonian in the following parts
$\tilde{H}=H_{\text{id}}+\kappa_{z}(\hat{n}+1/2)\hat{\sigma}_{z}+\kappa_{x}(2\hat{n}+1)\hat{\sigma}_{x}$ (39)
@@
$H_{\text{id}}$ $=\omega(\hat{a}^{\dagger}a+1/2)+\Omega^{(0)}\hat{\sigma}_{x},$ (40)
$\kappa_{z}$ $=2\delta^{(2)}l_{\text{ho}}^{2},$ (41)
$\kappa_{x}$ $=2\Omega^{(0)}(\delta^{(1)})^{2}l_{\text{ho}}^{2}/\omega^{2}+\Omega^{(2)}l_{\text{ho}}^{2}.$ (42)
We note that in Eq. (34), we can alternatively use the expression for the overlap of a squeezed-displaced Fock state from Ref. [45], rather than taking into account only the first few terms of the Baker-Hausdorff expansion. The overlap gives the transformed Rabi frequency for a given Fock state, $\Omega^{(0)}\innerproduct{-\alpha,-\xi,n}{\alpha,\xi,n}\hat{\sigma}_{x}$, which is well approximated by Eq. (34) when $\alpha$ and $\xi$ are small. The overlap also allows the calculation of off-resonant processes that change the Fock state $n\rightarrow m$, by $\Omega^{(0)}\innerproduct{-\alpha,-\xi,m}{\alpha,\xi,n}\hat{\sigma}_{x}$. This could be of interest, for instance to find tweezer parameters that specifically drive the sidebands $n\rightarrow n\pm 1$ or generate anharmonic potentials.

### v2

@@ V Author contributions
with quadrature transformations $\hat{D}^{\dagger}(\zeta\hat{\sigma}_{z})\hat{a}\hat{D}(\zeta\hat{\sigma}_{z})=\hat{a}-\zeta\hat{\sigma}_{z}$ and $\hat{D}^{\dagger}(\zeta\hat{\sigma}_{z})\hat{a}^{\dagger}\hat{D}(\zeta\hat{\sigma}_{z})=\hat{a}^{\dagger}-\zeta^{*}\hat{\sigma}_{z}$. The Hamiltonian becomes
$\tilde{H}=$ $\tilde{\omega}\left(\hat{a}^{\dagger}\hat{a}+1/2-\zeta\hat{\sigma}_{z}(\hat{a}+\hat{a}^{\dagger})+\zeta^{2}\right)$
$+\delta^{(0)}\hat{\sigma}_{z}+\delta^{(1)}e^{-{s\hat{\sigma}_{z}}}x_{0}(\hat{a}+\hat{a}^{\dagger})\hat{\sigma}_{z}$
$-2\delta^{(1)}e^{-{s\hat{\sigma}_{z}}}x_{0}\zeta$
$+\hat{D}^{\dagger}(\zeta\hat{\sigma}_{z})\hat{S}^{\dagger}(s\hat{\sigma}_{z})\hat{\Omega}\hat{\sigma}_{x}\hat{S}(s\hat{\sigma}_{z})\hat{D}(\zeta\hat{\sigma}_{z}),$ (S30)
@@
when $\zeta$ is real. The linear motion terms cancel out by choosing the displacement parameter $\zeta=\delta^{(1)}l_{\text{ho}}/\omega$ to be a dimensionless measure of the Stark shift gradient. The Hamiltonian simplifies to
$\tilde{H}=$ $\tilde{\omega}(\hat{a}^{\dagger}\hat{a}+1/2)-\tilde{\omega}\zeta^{2}+\delta^{(0)}\hat{\sigma}_{z}$
$+\hat{D}^{\dagger}(\zeta\hat{\sigma}_{z})\hat{S}^{\dagger}(s\hat{\sigma}_{z})\hat{\Omega}\hat{\sigma}_{x}\hat{S}(s\hat{\sigma}_{z})\hat{D}(\zeta\hat{\sigma}_{z}).$ (S31)
@@
Disregarding energy offsets that do not contribute to qubit-motion coupling, like $\delta^{(0)}\hat{\sigma}_{z}$ which can be fixed by the laser frequency, the Hamiltonian is
$\tilde{H}=$ $\tilde{\omega}(\hat{a}^{\dagger}\hat{a}+1/2)$
$+\hat{D}^{\dagger}(\zeta\hat{\sigma}_{z})\hat{S}^{\dagger}(s\hat{\sigma}_{z})\hat{\Omega}\hat{\sigma}_{x}\hat{S}(s\hat{\sigma}_{z})\hat{D}(\zeta\hat{\sigma}_{z}).$ (S32)
@@
The first term side describes the ion’s motion as a modified harmonic oscillator with frequency $\tilde{\omega}$. The second term describes the ion-laser coupling acted on by squeeze and displacement operators dependent on the qubit state of the ion. We evaluate the second term using the Baker-Hausdorff (BH) lemma, which gives corrections to the Rabi frequency amplitude $\Omega^{(0)}$. Disregarding off-resonant processes that change the Fock state,
$\hat{D}^{\dagger}(\zeta$ $\hat{\sigma}_{z})\hat{S}^{\dagger}(s\hat{\sigma}_{z})\left(\Omega^{(0)}\hat{\sigma}_{x}\right)\hat{S}(s\hat{\sigma}_{z})\hat{D}(\zeta\hat{\sigma}_{z})$ (S33)
$=\Omega^{(0)}\left(1+2\zeta^{2}(2\hat{n}+1)+\frac{s\zeta}{2}(2\hat{n}+1)+\order{\hat{n}^{2}}\right)\hat{\sigma}_{x}.$
The second term proportional to $(s\zeta/2)$ is suppressed by $10^{2}$ and thus negligible, along with higher order terms.
@@
We then calculate the squeeze and displacement transformation on the Rabi frequency curvature $\Omega^{(2)}\hat{x}^{2}$, which is expressed in creation and annihilation operators as $\Omega^{(2)}\hat{x}^{2}\hat{\sigma}_{x}=\Omega^{(2)}l_{\text{ho}}^{2}(2\hat{n}+1+\hat{a}^{\dagger}\hat{a}^{\dagger}+\hat{a}\hat{a})\hat{\sigma}_{x}$. Again using the BH lemma, we find
@@
$\hat{D}^{\dagger}(\zeta\hat{\sigma}_{z})\hat{S}^{\dagger}(s\hat{\sigma}_{z})\left(\Omega^{(2)}\hat{x}^{2}\hat{\sigma}_{x}\right)\hat{S}(s\hat{\sigma}_{z})\hat{D}(\zeta\hat{\sigma}_{z})$ $=\Omega^{(2)}l_{\text{ho}}^{2}(2\hat{n}+1)\left(1+2\zeta^{2}(2\hat{n}+1)+\frac{s\zeta}{2}(2\hat{n}+1)+\order{\hat{n}^{2}}\right)\hat{\sigma}_{x}$ (S34)
The second term proportional to $\zeta^{2}$ is suppressed by $\sim 10^{6}$, so we disregard it along with higher order terms. Inserting Eq. (S33) and Eq. (S34), the Hamiltonian is written in terms of the physical parameters as
@@
$\tilde{H}=$ $\omega(\hat{a}^{\dagger}\hat{a}+1/2)+\Omega^{(0)}\hat{\sigma}_{x}+2\delta^{(2)}l_{\text{ho}}^{2}(\hat{n}+1/2)\hat{\sigma}_{z}+\left((2\Omega^{(0)}(\delta^{(1)})^{2}l_{\text{ho}}^{2}/\omega^{2})(2\hat{n}+1)+\Omega^{(2)}l_{\text{ho}}^{2}(2\hat{n}+1)\right)\hat{\sigma}_{x}.$ (S35)
In the text this is written in the following parts
$\tilde{H}=H_{\text{id}}+\kappa_{z}(\hat{n}+1/2)\hat{\sigma}_{z}+\kappa_{x}(2\hat{n}+1)\hat{\sigma}_{x}$ (S36)
@@
$H_{\text{id}}$ $=\omega(\hat{a}^{\dagger}a+1/2)+\Omega^{(0)}\hat{\sigma}_{x},$ (S37)
$\kappa_{z}$ $=2\delta^{(2)}l_{\text{ho}}^{2},$ (S38)
$\kappa_{x}$ $=2\Omega^{(0)}(\delta^{(1)})^{2}l_{\text{ho}}^{2}/\omega^{2}+\Omega^{(2)}l_{\text{ho}}^{2}.$ (S39)
In our analytic model we have neglected off-resonant processes which change the Fock state $n$. In Fig. 3 we see that off-resonant terms become significant when $\bar{n}\ll 1$ or $\Omega\gg\omega$, by comparing the analytic and numerical curves. We note that instead of using the BH lemma we can utilize the expression for the overlap of a squeezed-displaced Fock state from Ref. [61]. The overlap gives the transformed Rabi frequency for a given Fock state, $\Omega^{(0)}\innerproduct{-\zeta,-s,n}{\zeta,s,n}\hat{\sigma}_{x}$, which is well approximated by the first few terms of the BH lemma when $\zeta$ and $s$ are small. The overlap also allows the calculation of off-resonant processes that change the Fock state $n\rightarrow m$, by $\Omega^{(0)}\innerproduct{-\zeta,-s,m}{\zeta,s,n}\hat{\sigma}_{x}$. This could be of interest, for instance to find tweezer parameters that specifically drive the sidebands $n\rightarrow n\pm 1$ or generate anharmonic potentials.
S3 Dependence on tweezer waist
We consider how the single qubit infidelity (Eq. (20)) varies with the beam waist $w_{0}$ in Fig. S6. As the waist increases non-paraxial errors become practically negligible. However, we note that increasing the waist is not always a valid solution in experimental setups - for instance when addressing multiple ions in a chain. Here, a larger waist can cause cross talk errors with adjacent ions [22]. In this case, one could find an optimal tweezer waist which minimizes both cross talk and non-paraxial effects. A more straightforward solution would be to use a magnetic field and polarization orientation that suppresses the non-paraxial error.
Figure S6: Analytic infidelity from Eq. 19 for the single qubit gate as a function of tweezer waist $w_{0}$, with different trap frequencies $\omega$. Here $\Omega^{(0)}=2\pi~{}0.2$ MHz and $\mathbb{B}=2.5$ G, and the ion temperature is fixed at (a) $\bar{n}=0.1$ and (b) $\bar{n}=1$.